OpenStudy (dumbsearch2):

Solve the following system using the substitution method: 2x = y 6x + 4y = -14

OpenStudy (aravindg):

ok giveb y=2x

OpenStudy (aravindg):

*given

OpenStudy (aravindg):

now replace y in second equation with 2x

OpenStudy (aravindg):

then solve for x

OpenStudy (aravindg):

having got x substitute its value in first equation to get y

Parth (parthkohli):

So you understand Aravind's explanation?

Parth (parthkohli):

LOL, I am red as long as I don't refresh

OpenStudy (dumbsearch2):

haha! ;D

OpenStudy (dumbsearch2):

no i don't understand. :(

OpenStudy (aravindg):

where exactly you have doubt @dumbsearch2 ?

OpenStudy (hba):

Plug in y=2x in eq(2)

OpenStudy (dumbsearch2):

how to find x

Parth (parthkohli):

\[\color{#C00}{2x} = y\]\[6x + 4\color{#T00}{\cdot y} = 6x + 4\cdot \color{#C00}{2x}\]

OpenStudy (dumbsearch2):

how to find x? cause to find y you need to know x right?

OpenStudy (hba):

how to find y ? cause to find x you need to know y right?

OpenStudy (aravindg):

you have an equation containing only x when you substitute y with x.Then solve for x

OpenStudy (aravindg):

gt it now @dumbsearch2 ?

OpenStudy (dumbsearch2):

nope. :( sorry I'm dumb...

OpenStudy (aravindg):

try to read the replies above again, try to understand what is said .

OpenStudy (callisto):

The first equation we have is 2x = y, that is the same as y=2x. The second equation we have is 6x + 4y = -14. Now, we put y=2x into the second equation. That is whenever we see "y" in the second equation, we re-write it as "2x" So, we get: 6x + 4 (2x) = -14 Can you solve this?

Parth (parthkohli):

You know that \(2x = y\), right? So can you say that \(6x + 4y =14\) can be written as \(6x + 4(2x) = 14\)?

OpenStudy (dumbsearch2):

x = -4 right?/

OpenStudy (dumbsearch2):

no i mean -1? no?

Parth (parthkohli):

Hmm... yes!

OpenStudy (dumbsearch2):

so y = -2! so y = -2, and x= -1! no?

Parth (parthkohli):

Sorry, it's not \(6x + 4y = 14\). I got my reply wrong :-P

Parth (parthkohli):

Yesh!

OpenStudy (callisto):

*Clapping hands*

Parth (parthkohli):

Yay!

OpenStudy (dumbsearch2):

:DD thanks so MUCHJ GUYS! im trying so hard to finish math online course for tomrrow... it closes on tomorrow, and I've got 2 units to do online!!! (each unit has like 4 lessons and 4 quizes) I'm kinda bad at the subtituation method and stuff, wish I would understand! it techinically is tomorrow, I mean it's the very very very early morning right now!!i know! I'm trying to use openstudy to help me too... (that's how I stumbled upon this site today!)

OpenStudy (dumbsearch2):

So like this! (Confirm! :))

Parth (parthkohli):

Yes :-)

OpenStudy (dumbsearch2):

yeaa! ;D

OpenStudy (dumbsearch2):

got a 75% grade in the quiz, not bad. :D